a. Now let’s go through the same financial exercise using a savings of $90 a month or $1080 a year (about $3 a day). While your current savings (or starting balance) is $0, you create an automatic deposit of $90 a month starting with your first paycheck. In other words, assume you directly deposit $90 a month into a well diversified investment account earning 7% interest compounded yearly from now until you retire at age 67. How much would you have in the account 45 years from now when you retire at 67? (Enter numbers only, no dollar sign or comma. For example, if your answer is $275,421 then enter 275421 as your submission.) 4________

(b). ​The total value from Question (a)  is divided into two parts. How much is from total interest? (Enter numbers only, no dollar sign or comma.) ______

(c) Compare the total interest earned from saving $90 per month to the total interest earned from saving $60 per month. How much more interest have you accumulated by age 67 from saving $90 per month? ​ (Enter numbers only, no dollar sign or comma.) $____________
Answer #1

I = 7% = 7% / 12 = 0.58333% per month

t = 45 yrs = 45 * 12 = 540

A =90

F = 90 * (F/A,0.58333%,540) = 90* 3792.5454 = 341329.09 = 341329 (nearest dollar)

Interest = 341329.09 - 90*540 = 292729 (nearest dollar)

When A = 60

F = 60 * (F/A,0.58333%,540) = 60* 3792.5454 = 227552.72

interest = 227552.72 - 60*540 = 195152.72 = 195153

difference in Interest earned = 292729 - 195153 = 97576